Write the successor of:
(a) 2440701
(b) 100199
(c) 1099999
 
egbjmu89erweghi84J9U85HUJITEJMH8I5TJ

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

The successors are

(a) 2440701 + 1 = 2440702

(b) 100199 + 1 = 100200

(c) 1099999 + 1 = 1100000


Thanks